Nick runs around the perimeter of a circular track at 11 ft/sec. The track has a radius of 99 yd. After 30 seconds, Nick turns and runs to the center of the track along a straight (radial) line. Upon reaching the center, he turns and runs away along a different radial line to his starting position on the perimeter.

How far has Nick traveled once he has returned to his starting position?

distance =  feet


1
Expert's answer
2021-01-26T04:43:32-0500


Here,

we have a circle centered at O and A is the starting point of Nick, with OA as the radius of the circle.

We know OA=99yd=297ft.

Then Nick runs for 30 seconds till B on the circumference and then returns to O and then back to A via

a radial line.



"OA=99yd=297ft"

Now, Nick is running at a speed of "11ft\/sec" for "30" seconds.

So, Nick travelled "11\cdot30 ft=330ft" for those "30" seconds. (as distance= speed"\cdot" time)


"\therefore" "\overgroup{AB}=330ft"


Now, he goes to the center via OB="297ft" and then further to his starting point A via OA="297ft" .


So, the total distance covered by Nick is


"\overgroup{AB}+\overline{BO}+\overline{OA}\\
=330+297+297=924ft"

Therefore the total distance Nick travelled is "924ft" .